# Parc Científic UV, c/ Catedrático José Beltrán, 2, E-46980 Paterna (València), Spain

### v2

@@ 3.3 Diagonal plus off-diagonal mass matrix
3.3.1 Tree level: the Linear plus Inverse Seesaw Model
@@
to the Lagrangian of the LSS model, $\mathcal{L}_{\rm LSS}$ given in Eq. (40), [^8] One can also add another one in the $M_{\nu}^{11}$ element, but it will only contribute to neutrino masses at one loop. See Ref. Gavela:2009cd ; Dev:2012sg ; CentellesChulia:2020dfh for other variants of the inverse seesaw. so that:
@@
where the first (second) contribution corresponds to the linear (inverse) seesaw. This has the form of Eq. (3) with
$Y_{1}=Y_{S}^{T}\,,\qquad Y_{2}=Y_{N}\,,\qquad M_{22}=\frac{v^{2}}{2}{\mathrm{M}_{R}^{T}}^{-1}\,\mu\,{\mathrm{M}_{R}^{-1}}\,,\qquad M_{12}=\frac{v^{2}}{2}\mathrm{M}_{R}^{-1}\,.$ (57)
@@
$V^{\ast}\begin{pmatrix}0&M_{12}\\ M_{12}&M_{22}\end{pmatrix}V^{\dagger}\approx\begin{pmatrix}M_{12}+M_{22}/2&0\\ 0&M_{12}-M_{22}/2\end{pmatrix}\,$, (58)
@@
${Y_{1}}=$ $\dfrac{e^{-\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\mp i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]+\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\dfrac{M_{22}}{4M_{12}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$, (59)
$Y_{2}=$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$,
@@
${Y_{1}}=$ $\dfrac{e^{-\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\mp i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]+\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\dfrac{M_{22}}{4M_{12}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$, (60)
$Y_{2}=$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$,
where $u_{i}$ are the two columns of the PMNS mixing matrix. For completeness, we remark that also the $Y_{2}$ Yukawa receive contributions proportional to the LNV term; however, these are further suppressed by a factor $e^{-\text{Imz}}$ and are thus neglected. As a consistency check, we find that our results reproduce those of Ref. Hernandez:2022ivz upon identifying $M_{22}\rightarrow\Delta M$ and $M_{12}\rightarrow M$.
3.3.2 Loop level: the Extended Scotogenic Model
@@
Much like in the periodic table, where the presence of gaps has guided the search for missing elements, the absence of a radiative model incorporating both diagonal and off-diagonal mass contributions motivates the following proposal. The model is the generalisation of the Scotogenic Model (ScM) Tao96 ; Ma:2006km , still based on a $\mathbb{Z}_{2}$ symmetry, but adding both chiralities of the fermion singlet. We term it the Extended Scotogenic Model (EScM). The particle content with the associated charges can be found in Table 3. The Lagrangian for the model is:
$\mathcal{L}_{\rm EScM}=\bar{\psi}m_{D}\psi+\frac{1}{2}\overline{\psi_{L}^{c}}m_{L}\psi_{L}+\frac{1}{2}\overline{\psi_{R}^{c}}m_{R}\psi_{R}+\left(\overline{\psi_{R}}\tilde{\Phi}^{\dagger}YL+\overline{\psi_{L}}\tilde{\Phi}^{T}Y^{\prime*}\tilde{L}+\text{H.c.}\right),$ (61)
where for the case of just one copy of the fermion $\psi$, the neutrino Yukawa couplings $Y,Y^{\prime}$ are three-component row vectors. The scalar potential is provided in Eq. (24). In the basis $(\nu_{L},\psi_{R}^{c},\psi_{L})$, the complete neutral fermion mass matrix at tree-level reads:
@@
Light neutrino masses are generated at one loop, as shown in the diagrams of Fig. 4. The neutrino mass matrix can be written as in Eq. (4) with [^9] Note that the contributions to the neutrino mass matrix can be equivalently expressed in the form of Eq. (3).
$\mathcal{Y}=\begin{pmatrix}Y\\ Y^{\prime}\end{pmatrix}\,,\quad\mathcal{M}=\frac{1}{32\pi^{2}}\begin{pmatrix}F_{1}M_{1}s_{\theta}^{2}+F_{2}M_{2}c_{\theta}^{2}&(F_{1}M_{1}-F_{2}M_{2})s_{\theta}c_{\theta}\\ (F_{1}M_{1}-F_{2}M_{2})s_{\theta}c_{\theta}&F_{1}M_{1}c_{\theta}^{2}+F_{2}M_{2}s_{\theta}^{2}\end{pmatrix}\,$, (68)
where $s_{\theta}=\sin\theta$ and $c_{\theta}=\cos\theta$, with the mixing angle $\theta$ defined in Eq. (65). Here $F_{k}\equiv\mathcal{F}(m_{\phi_{0}^{R}},m_{\phi_{0}^{I}},M_{k})$, where the loop function $\mathcal{F}$ is given in Eq. (29), and $M_{1}$ and $M_{2}$ are the mass eigenstates in Eq. (63). Importantly, even for the case of only one family of heavy fermions, two masses for light neutrinos are generated. It is therefore possible to use the GCI parametrisation, see also Ref. Avila:2019hhv . First, let us consider the seesaw case for the heavy fermions, $m_{R}\gg m_{D}\gg m_{L}$, such that $s_{\theta}\simeq m_{D}/m_{R}$. The elements of $\mathcal{M}$ are given by:
$M_{11}\simeq\frac{m_{R}}{32\pi^{2}}F_{2}\,,\quad M_{12}\simeq-\frac{m_{D}}{32\pi^{2}}F_{2}\,,\quad M_{22}\simeq\frac{1}{32\pi^{2}}\left[m_{L}\,F_{1}+(F_{2}-F_{1})\frac{m^{2}_{D}}{m_{R}}\right]\,.$ (69)
@@
In this limit, the conventional seesaw scenario is reproduced, allowing to directly use the GCI parametrisation. On the other hand, expanding $\mathcal{M}$ in Eq. (68) in the pseudo-Dirac limit, $m_{D}\gg m_{L,R}$, we have $s_{\theta}\simeq 1/\sqrt{2}$, and obtain:
@@
$m_{\nu}=\frac{1}{32\pi^{2}}(F_{1}+F_{2})\left[Y^{T}\,m_{R}\,Y+Y^{\prime T}\,m_{L}\,Y^{\prime}+Y^{T}\,m_{D}\,Y^{\prime}+Y^{\prime T}\,m_{D}\,Y\right]\,.$ (73)
We thus recover the same neutrino mass structure discussed in the LSS+ISS case, with the difference being that in the present scenario $M_{11}\not=0$. The presence of this term modifies the form of the eigenvalues of $\mathcal{M}$, which now receive an additional contribution proportional to it, leading to an enhanced splitting between the eigenvalues of $\mathcal{M}$ (as discussed in Ref. Hernandez:2022ivz for the LSS + ISS case). The parametrisation of the Yukawa couplings, however, remains unchanged in its overall structure, and given by Eqs. (59) and (60) with $M_{22}$ now replaced by $M_{11}+M_{22}$.
4 Implementation of the Generalised Casas-Ibarra parametrisation in models with extra symmetries
@@ 4.1 The Zee Model
As an illustrative example, we consider the Zee model Zee:1980ai ; Cheng:1980qt ; Wolfenstein:1980sy , see also Ref. Herrero-Garcia:2017xdu for a full analysis. The model provides an example of one-loop neutrino mass generation and consists of an extension of the Standard Model by an additional scalar doublet $H_{2}$, together with a singly-charged scalar singlet $h^{+}$. The particle content of the model is summarised in Table 4 and the neutrino mass generation at one-loop is shown in Fig. 5.
@@
The explicit form of the matrix $V$ in $D_{\mathcal{M}}=V^{\ast}\mathcal{M}V^{\dagger}$ is given in Eq. (33). This allows for the use of the GCI parametrisation, Eq. (11). However, in this scenario, the semi-orthogonal complex matrix $\mathcal{R}$ is not arbitrary, but must be such that it realises the antisymmetric structure of the Yukawa couplings associated with the singly-charged scalar $h^{+}$, i.e., $f^{T}=-f$.
@@
Then it is possible to choose 6 free (complex) parameters and obtain analytical expressions for the remaining ones in the following way. Two of the three Yukawa couplings in $\hat{f}$ can be freely chosen, while the third one is fixed by the condition that $\hat{f}$, being of rank 2, has a null eigenvector Felkl:2021qdn . Specifically, writing $\hat{f}$ as
@@
leading to a quadratic equation that fixes one of the elements of $\hat{f}$ in terms of the other two Felkl:2021qdn . Moreover, once $\hat{f}$ is fixed, the matrix $A$ is completely determined (for a given light neutrino mass matrix) and therefore Eq. (86) leads to a linear system of 6 equations for the 9 elements of $R_{2}$. Actually, given that $\hat{f}$ is not invertible, neither is $A$, and one can show that there are only five independent linear equations for the elements of $R_{2}$. The linear system can always be solved analytically, yielding 5 elements of $R_{2}$ in terms of the other four that are chosen as free parameters. In turn, $R_{1}$ is determined from Eq. (84) (which can also be used to choose a different set of free parameters).
@@
$\begin{pmatrix}r_{23}\\[4.0pt] r_{33}\end{pmatrix}$ $=\frac{1}{\lambda_{2}\,d}\begin{pmatrix}a_{33}\quad&-(\lambda_{2}a_{32}+\lambda_{3}a_{33})\\[4.0pt] -a_{23}\quad&\lambda_{2}a_{22}+\lambda_{3}a_{23}\end{pmatrix}\,\begin{pmatrix}b_{13}-\sum_{j=1}^{3}a_{j3}\,r_{j1}-\left(\lambda_{2}\,a_{12}+\lambda_{3}\,a_{13}\right)\,r_{13}\\[4.0pt] b_{33}/2-a_{13}\,r_{13}\end{pmatrix}\,$,
$\begin{pmatrix}r_{22}\\[4.0pt] r_{32}\end{pmatrix}$ $=\frac{1}{d}\begin{pmatrix}a_{33}\quad&-a_{32}\\[4.0pt] -a_{23}\quad&a_{22}\end{pmatrix}\,\begin{pmatrix}b_{22}/2-a_{12}\,r_{12}\\[4.0pt] b_{23}-a_{13}\,r_{12}-\sum_{j=1}^{3}a_{j2}\,r_{j3}\end{pmatrix}\,$, (95)
@@
This procedure gives a parametrisation of the Zee model and other models with an antisymmetric Yukawa matrix, with analytical expressions for all the parameters in terms of a chosen set of free parameters. A detailed application of this method is left for future work. We end up noting that, once the Yukawa couplings in $f$ are fixed, instead of following the GCI procedure outlined above, it would also be possible to set and solve analytically a linear system of equations directly for the Yukawa couplings in $\Gamma_{e}$. This approach and the comparison to the one outlined above are also left for future work.
5 The complete mass matrix with right-handed neutrinos: Discussion and Outlook

### v3

@@ III.3 Diagonal plus off-diagonal mass matrix
III.3.1 Tree level: the Linear plus Inverse Seesaw Model
@@
to the Lagrangian of the LSS model, $\mathcal{L}_{\rm LSS}$ given in Eq. (38), [^8] One can also add another one in the $M_{\nu}^{11}$ element, but it will only contribute to neutrino masses at one loop. See Ref. [45, 37, 26] for other variants of the inverse seesaw. so that:
@@
where the first (second) contribution corresponds to the linear (inverse) seesaw. This has the form of Eq. (II) with
$Y_{1}=Y_{S}^{T},\,Y_{2}=Y_{N},\,$
$M_{22}=\frac{v^{2}}{2}{\mathrm{M}_{R}^{T}}^{-1}\,\mu\,{\mathrm{M}_{R}^{-1}},\,M_{12}=\frac{v^{2}}{2}\mathrm{M}_{R}^{-1}\,.$ (55)
@@
$V^{\ast}\begin{pmatrix}0&M_{12}\\ M_{12}&M_{22}\end{pmatrix}V^{\dagger}\approx\begin{pmatrix}M_{12}+\frac{M_{22}}{2}&0\\ 0&M_{12}-\frac{M_{22}}{2}\end{pmatrix}\,$, (56)
@@
${Y_{1}}=$ $\dfrac{e^{-\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\mp i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]$
$+$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\dfrac{M_{22}}{4M_{12}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$,
$Y_{2}=$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{1}^{\ast}\sqrt{m_{1}}\pm i\,u_{2}^{\ast}\sqrt{m_{2}}\Big]\,$, (57)
@@
${Y_{1}}=$ $\dfrac{e^{-\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\mp i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]$
$+$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\dfrac{M_{22}}{4M_{12}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$,
$Y_{2}=$ $\dfrac{e^{\text{Im}z}}{\sqrt{2M_{12}}}\Big[u_{2}^{\ast}\sqrt{m_{2}}\pm i\,u_{3}^{\ast}\sqrt{m_{3}}\Big]\,$, (58)
where $u_{i}$ are the two columns of the PMNS mixing matrix. For completeness, we remark that also the $Y_{2}$ Yukawa receive contributions proportional to the LNV term; however, these are further suppressed by a factor $e^{-\text{Imz}}$ and are thus neglected. As a consistency check, we find that our results reproduce those of Ref. [57] upon identifying $M_{22}\rightarrow\Delta M$ and $M_{12}\rightarrow M$.
III.3.2 Loop level: the Extended Scotogenic Model
@@
Much like in the periodic table, where the presence of gaps has guided the search for missing elements, the absence of a radiative model incorporating both diagonal and off-diagonal mass contributions motivates the following proposal. The model is the generalisation of the Scotogenic Model (ScM) [93, 72], still based on a $\mathbb{Z}_{2}$ symmetry, but adding both chiralities of the fermion singlet. We term it the Extended Scotogenic Model (EScM). The particle content with the associated charges can be found in Table 3. The Lagrangian for the model is:
$-\mathcal{L}_{\rm EScM}$ $=\bar{\psi}m_{D}\psi+\frac{1}{2}\overline{\psi_{L}^{c}}m_{L}\psi_{L}+\frac{1}{2}\overline{\psi_{R}^{c}}m_{R}\psi_{R}$
$+\left(\overline{\psi_{R}}\tilde{\Phi}^{\dagger}YL+\overline{\psi_{L}}\tilde{\Phi}^{T}Y^{\prime*}\tilde{L}+\text{H.c.}\right),$ (59)
where for the case of just one copy of the fermion $\psi$, the neutrino Yukawa couplings $Y,Y^{\prime}$ are three-component row vectors. The scalar potential is provided in Eq. (III.1.2). In the basis $(\nu_{L},\psi_{R}^{c},\psi_{L})$, the complete neutral fermion mass matrix at tree-level reads:
@@
Light neutrino masses are generated at one loop, as shown in the diagrams of Fig. 4. The neutrino mass matrix can be written as in Eq. (4) with [^9] Note that the contributions to the neutrino mass matrix can be equivalently expressed in the form of Eq. (II).
$\mathcal{Y}=\begin{pmatrix}Y\\ Y^{\prime}\end{pmatrix}\,$,
$\mathcal{M}=\frac{1}{32\pi^{2}}\begin{pmatrix}F_{1}M_{1}s_{\theta}^{2}+F_{2}M_{2}c_{\theta}^{2}&(F_{1}M_{1}-F_{2}M_{2})s_{\theta}c_{\theta}\\ (F_{1}M_{1}-F_{2}M_{2})s_{\theta}c_{\theta}&F_{1}M_{1}c_{\theta}^{2}+F_{2}M_{2}s_{\theta}^{2}\end{pmatrix}\,$, (66)
where $s_{\theta}=\sin\theta$ and $c_{\theta}=\cos\theta$, with the mixing angle $\theta$ defined in Eq. (63). Here $F_{k}\equiv\mathcal{F}(m_{\phi_{0}^{R}},m_{\phi_{0}^{I}},M_{k})$, where the loop function $\mathcal{F}$ is given in Eq. (27), and $M_{1}$ and $M_{2}$ are the mass eigenstates in Eq. (61). Importantly, even for the case of only one family of heavy fermions, two masses for light neutrinos are generated. It is therefore possible to use the GCI parametrisation, see also Ref. [8]. First, let us consider the seesaw case for the heavy fermions, $m_{R}\gg m_{D}\gg m_{L}$, such that $s_{\theta}\simeq m_{D}/m_{R}$. The elements of $\mathcal{M}$ are given by:
$M_{11}\simeq\frac{m_{R}}{32\pi^{2}}F_{2}\,$,
$M_{12}\simeq-\frac{m_{D}}{32\pi^{2}}F_{2}\,$,
$M_{22}\simeq\frac{1}{32\pi^{2}}\left[m_{L}\,F_{1}+(F_{2}-F_{1})\frac{m^{2}_{D}}{m_{R}}\right]\,.$ (67)
@@
In this limit, the conventional seesaw scenario is reproduced, allowing to directly use the GCI parametrisation. On the other hand, expanding $\mathcal{M}$ in Eq. (III.3.2) in the pseudo-Dirac limit, $m_{D}\gg m_{L,R}$, we have $s_{\theta}\simeq 1/\sqrt{2}$, and obtain:
@@
$m_{\nu}$ $=\frac{1}{32\pi^{2}}(F_{1}+F_{2})[Y^{T}\,m_{R}\,Y+Y^{\prime T}\,m_{L}\,Y^{\prime}$
$+Y^{T}\,m_{D}\,Y^{\prime}+Y^{\prime T}\,m_{D}\,Y]\,.$ (71)
We thus recover the same neutrino mass structure discussed in the LSS+ISS case, with the difference being that in the present scenario $M_{11}\not=0$. The presence of this term modifies the form of the eigenvalues of $\mathcal{M}$, which now receive an additional contribution proportional to it, leading to an enhanced splitting between the eigenvalues of $\mathcal{M}$ (as discussed in Ref. [57] for the LSS + ISS case). The parametrisation of the Yukawa couplings, however, remains unchanged in its overall structure, and given by Eqs. (III.3.1) and (III.3.1) with $M_{22}$ now replaced by $M_{11}+M_{22}$.
IV Implementation of the Generalised Casas-Ibarra parametrisation in models with extra symmetries
@@ IV.1 The Zee Model
As an illustrative example, we consider the Zee model [102, 30, 95], see also Ref. [60] for a full analysis. The model provides an example of one-loop neutrino mass generation and consists of an extension of the Standard Model by an additional scalar doublet $H_{2}$, together with a singly-charged scalar singlet $h^{+}$. The particle content of the model is summarised in Table 4 and the neutrino mass generation at one-loop is shown in Fig. 5.
@@
The explicit form of the matrix $V$ in $D_{\mathcal{M}}=V^{\ast}\mathcal{M}V^{\dagger}$ is given in Eq. (31). This allows for the use of the GCI parametrisation, Eq. (11). However, in this scenario, the matrix $\mathcal{R}$ has less degrees of freedom than an arbitrary semi-orthogonal complex matrix, because it must realize the antisymmetric structure of the Yukawa matrix associated with the singly-charged scalar $h^{+}$, i.e., $f^{T}=-f$.
@@
Then it is possible to choose 6 free (complex) parameters and obtain analytical expressions for the remaining ones in the following way. Two of the three Yukawa couplings in $\hat{f}$ can be freely chosen, while the third one is fixed by the condition that $\hat{f}$, being of rank 2, has a null eigenvector [42]. Specifically, writing $\hat{f}$ as
@@
leading to a quadratic equation that fixes one of the elements of $\hat{f}$ in terms of the other two [42]. Moreover, once $\hat{f}$ is fixed, the matrix $A$ is completely determined (for a given light neutrino mass matrix) and therefore Eq. (84) leads to a linear system of 6 equations for the 9 elements of $R_{2}$. Actually, given that $\hat{f}$ is not invertible, neither is $A$, and one can show that there are only five independent linear equations for the elements of $R_{2}$. The linear system can always be solved analytically, yielding 5 elements of $R_{2}$ in terms of the other four that are chosen as free parameters. In turn, $R_{1}$ is determined from Eq. (82) (which can also be used to choose a different set of free parameters).
@@
$\begin{pmatrix}r_{23}\\ r_{33}\end{pmatrix}$ $=\frac{1}{\lambda_{2}\,d}\begin{pmatrix}a_{33}\quad&-(\lambda_{2}a_{32}+\lambda_{3}a_{33})\\[4.0pt] -a_{23}\quad&\lambda_{2}a_{22}+\lambda_{3}a_{23}\end{pmatrix}\,\times$
$\begin{pmatrix}b_{13}-\sum_{j=1}^{3}a_{j3}\,r_{j1}-\left(\lambda_{2}\,a_{12}+\lambda_{3}\,a_{13}\right)\,r_{13}\\[4.0pt] b_{33}/2-a_{13}\,r_{13}\end{pmatrix}\,$,
$\begin{pmatrix}r_{22}\\[4.0pt] r_{32}\end{pmatrix}$ $=\frac{1}{d}\begin{pmatrix}a_{33}&-a_{32}\\ -a_{23}&a_{22}\end{pmatrix}\,\begin{pmatrix}b_{22}/2-a_{12}\,r_{12}\\ b_{23}-a_{13}\,r_{12}-\sum_{j=1}^{3}a_{j2}\,r_{j3}\end{pmatrix}\,$, (93)
@@
This procedure gives a parametrisation of the Zee model and other models with an antisymmetric Yukawa matrix, with analytical expressions for all the parameters in terms of a chosen set of free parameters. Note that once the Yukawa couplings in $f$ are fixed, instead of following the GCI procedure outlined above, it would also be possible to set and solve analytically a linear system of equations directly for the Yukawa couplings in $\Gamma_{e}$. Also notice that a different parametrisation was proposed in Ref. [73] [^11] Note that in the Zee-Wolfenstein model [95] only one scalar doublet couples to the SM leptons (imposed via a discrete symmetry), and therefore $\rm{Tr}\,(m_{\nu})=0$. However, this possibility is now excluded by data. Therefore, in general, both scalar doublets couple to the leptons and $\rm{Tr}\,(m_{\nu})\neq 0$. and it was applied to phenomenological studies in Refs. [55, 36]. Further insight into the relationship among the different parameters of the Zee model and observational constraints, particularly on charged lepton flavour violation, was given in Ref. [29]. Since different parametrisations may have different virtues, comparing and applying these approaches could be an interesting direction for future work.
V The complete mass matrix with right-handed neutrinos: Discussion and Outlook